# Truncated-degree-choosability of planar graphs

### v1

@@ Abstract
Assume $G$ is a graph and $k$ is a positive integer. Let $f:V(G)\to\mathbb{N}$ be defined as $f(v)=\min\{k,d_{G}(v)\}$. If $G$ is $f$-choosable, then we say $G$ is $k$-truncated-degree-choosable. It was proved in [Zhou,Zhu,Zhu, Arc-weighted acyclic orientations and variations of degeneracy of graphs, arXiv:2308.15853] that there is a 3-connected non-complete planar graph that is not 7-truncated-degree-choosable, and every 3-connected non-complete planar graph is 16-truncated-degree-choosable. This paper improves the bounds, and proves that there is a 3-connected non-complete planar graph that is not 8-truncated-degree-choosable and every 3-connected non-complete planar graph is 12-truncated-degree-choosable.
Keywords Degree-choosable, Truncated-degree-choosable, Planar graphs
@@ 1 Introduction
A list assignment $L$ for a graph $G$ is a mapping that assigns a set $L(v)$ of permissible colours to each vertex $v$ of $G$. We denote by $\mathbb{N}^{G}$ the set of mappings from $V(G)$ to =$\mathbb{N}={\left\{0,1,\ldots\right\}}$. For $f\in\mathbb{N}^{G}$, an $f$-list assignment is an assignment $L$ with $|L(v)|\geq f(v)$ for each vertex $v$. An $L$-colouring of $G$ is a mapping $\phi$ that assigns a colour $\phi(v)$ from $L(v)$ to each vertex $v$, such that $\phi(u)\neq\phi(v)$ for every edge $uv$ of $G$. We say $G$ is $f$-choosable if $G$ is $L$-colourable for any $f$-list assignment $L$ of $G$. The choice number $\operatorname{ch}(G)$ of $G$ is the minimum integer $k$ such that $G$ is $k$-choosable (i.e., $f$-choosable for the constant mapping $f\in\mathbb{N}^{G}$ defined as $f(v)=k$ for all $v$). The choice number is an important graph invariant that has been studied extensively in the literature (cf. [5]).
A graph $G$ is called degree-choosable if it is $f$-choosable for the mapping $f\in\mathbb{N}^{G}$ defined as $f(v)=d_{G}(v)$ for $v\in V(G)$, where $d_{G}(v)$ is the degree of vertex $v$. Degree-choosable graphs are characterized in [3] and [7]: A connected graph $G$ is not degree-choosable if and only if $G$ is a Gallai-tree, i.e., each block of $G$ is either a complete graph or an odd cycle.
In this paper, we investigate a combination of $k$-choosability and degree-choosability, which is called $k$-truncated-degree-choosability (cf. [8]).
@@
Assume $G$ is a graph and $k$ is a positive integer. Let $f\in\mathbb{N}^{G}$ be defined as $f(v)=\min\{k,d_{G}(v)\}$ for each vertex $v\in V(G)$. If $G$ is $f$-choosable, then we say $G$ is $k$-truncated-degree-choosable.
Truncated-degree-choosability of graphs was first studied by Hutchinson [4]. Bruce Richter asked whether every 3-connected non-complete planar graph is 6-truncated-degree-choosable. Note that for any positive integer $k$, the complete bipartite graph $K_{2,k^{2}}$ is planar and not $k$-truncated-degree-choosable. Also complete graphs are not degree-choosable. So for the study of $k$-truncated choosability of planar graphs, the restriction to 3-connected non-complete graphs is natural. Motivated by Richter’s question, Hutchinson studied truncated-degree-choosability of outerplanar graphs. It was proved in [4] that if $G\neq K_{3}$ is a 2-connected maximal outerplanar graphs (i.e., each inner face is a triangle), then $G$ is 5-truncated-degree-choosable. If $G$ is a 2-connected bipartite outerplanar graphs, then $G$ is 4-truncated-degree-choosable. The results are sharp, as there are 2-connected maximal outerplanar graphs that are not 4-truncated-degree-choosable, and 2-connected bipartite outplanar graphs that are not 3-truncated-degree-choosable. Hutchinson’s result is strengthened in [2], where it is shown that 2-connected outerplanar graphs other than odd cycles are 5-truncated-degree-DP-colourable, and hence 5-truncated-degree-choosable. This result is further strengthened in [6], where it is proved that 2-connected $K_{2,4}$-minor free graphs other than cycles and complete graphs are 5-truncated-degree-DP-colourable. Truncated-degree-choosability of graphs were also studied in [1]. Some families of graphs are shown to be $k$-truncated-degree-choosable, by putting restrictions on the distance between components of the subgraph induced by vertices of degree less than $k$.
Recently, Zhou, Zhu, and Zhu [8] answered Richter’s question in negative and constructed a 3-connected non-complete planar graph that is not 7-truncated-degree-choosable. Then they proved that every 3-connected non-complete planar graph is $ST^{(2)}$-$16$-truncated-degree-degenerate, which implies that these graphs are 16-truncated-degree-choosable.
@@
Every 3-connected non-complete planar graph is 12-truncated-degree-choosable.
@@
There is a 3-connected non-complete planar graph which is not 8-truncated-degree-choosable.
We remark that the result in [8] implies that 3-connected non-complete planar graphs are 16-truncated-degree-DP-paintable and 16-truncated-degree-AT. Our result takes advantage of the list assignment, and does not apply to DP-colouring, on-line list colouring and Alon-Tarsi orientation.
2 Some preliminaries
The following lemmas proved in [1] and [8] are needed in our proof. If $G$ is a Gallai-tree, and $L$ is a list assignment of $G$ with $|L(v)|\geq d_{G}(v)$ such that $G$ is not $L$-colourable, then we say $L$ is a bad list assignment for $G$.
Lemma 1 ([1]).
Note that each block of a Gallai-tree is a regular graph. It follows from Lemma 1 that for a bad list assignment $L$ of a Gallai-tree $G$, $|L(v)|=d_{G}(v)$ for each vertex $v\in V(G)$. One property of a bad list assignment $L$ we shall use frequently is that if two vertices $u,v$ are in a same block of $G$ and none of them is a cut-vertex of $G$, then $L(u)=L(v)$.
Assume $G$ is a simple plane graph. We denote by $F(G)$ the set of faces of $G$. For each face $\theta$ of $G$, let $V(\theta)$ be the set of vertices on the boundary of $\theta$. Let $\Theta(G)$ be the bipartite graph with partite sets $V(G)$ and $F(G)$, where $v\theta$ is an edge in $\Theta(G)$ if and only if $v\in V(\theta)$.
Note that $G$ need not be connected. Assume $G$ has connected components $G_{1},G_{2},\ldots,G_{k}$. For a face $\theta$ of $G$, let $J_{\theta}=\{i:V(\theta)\cap V(G_{i})\neq\emptyset\}$. For $i\in J_{\theta}$, let $\theta_{i}$ be the face of $G_{i}$ that contains $\theta$. Then $\theta=\cap_{i\in J_{\theta}}\theta_{i}$ and $V(\theta)=\cup_{i\in J_{\theta}}V\left(\theta_{i}\right)$.
Definition 2.
Let $\theta^{*}$ be the infinite face of $G$, and $v^{*}$ be an arbitrary vertex in $V(\theta^{*})$. A spanning subgraph $F$ of $\Theta(G)$ is very nice with respect to $(\theta^{*},v^{*})$ if the following hold:
  • $d_{F}(v^{*})=1$ and $d_{F}(v)\leq 2$ for $v\in V(G)-\{v^{*}\}$.
  • If $v_{1},v_{2}\in V(\theta)-N_{F}(\theta)$, then $v_{1},v_{2}$ are contained in a cycle $C$ of $V(\theta)$.
Lemma 2 ([8]).
For any simple plane graph $G$, $\Theta(G)$ has a very nice subgraph.

### v2

# Degree-truncated choice number of planar graphs

@@ Abstract
Assume $G$ is a graph and $k$ is a positive integer. Let $f:V(G)\to\mathbb{N}$ be defined as $f(v)=\min\{k,d_{G}(v)\}$. If $G$ is $f$-choosable, then we say $G$ is degree-truncated $k$-choosable. Answering a question of Richter, it was proved in [Zhou,Zhu,Zhu, Degree-truncated choice number of graphs, arXiv:2308.15853] that there exists a 3-connected non-complete planar graph that is not degree-truncated 7-choosable, and every 3-connected non-complete planar graph is degree-truncated 16-choosable. This paper improves the bounds, and proves that there exists a 3-connected non-complete planar graph that is not degree-truncated 8-choosable, and that every 3-connected non-complete planar graph is degree-truncated $12$-choosable.
Keywords List colouring,  choice number,  degree-choosable, degree-truncated $k$-choosable, planar graphs
@@ 1 Introduction
A list assignment $L$ for a graph $G$ is a mapping that assigns a set $L(v)$ of permissible colours to each vertex $v$ of $G$. We denote by $\mathbb{N}^{G}$ the set of mappings from $V(G)$ to =$\mathbb{N}={\left\{0,1,\ldots\right\}}$. For $f\in\mathbb{N}^{G}$, an $f$-list assignment is an assignment $L$ with $|L(v)|\geq f(v)$ for each vertex $v$. An $L$-colouring of $G$ is a mapping $\phi$ that assigns a colour $\phi(v)$ from $L(v)$ to each vertex $v$, such that $\phi(u)\neq\phi(v)$ for every edge $uv$ of $G$. We say $G$ is $f$-choosable if $G$ is $L$-colourable for any $f$-list assignment $L$ of $G$, and say $G$ is $k$-choosable if $G$ is $f$-choosable for the constant mapping $f\in\mathbb{N}^{G}$ defined as $f(v)=k$ for all $v$. The choice number $\operatorname{ch}(G)$ of $G$ is the minimum integer $k$ such that $G$ is $k$-choosable. We say $G$ is degree-choosable if it is $f$-choosable for the mapping $f\in\mathbb{N}^{G}$ defined as $f(v)=d_{G}(v)$ for $v\in V(G)$, where $d_{G}(v)$ is the degree of vertex $v$. Degree-choosable graphs are characterized in [2] and [8]: A connected graph $G$ is not degree-choosable if and only if $G$ is a Gallai-tree, i.e., each block of $G$ is either a complete graph or an odd cycle.
In this paper, we investigate a combination of $k$-choosability and degree-choosability, which is called degree-truncated $k$-choosability in [9].
@@
Assume $G$ is a graph and $k$ is a positive integer. Let $f\in\mathbb{N}^{G}$ be defined as $f(v)=\min\{k,d_{G}(v)\}$ for each vertex $v\in V(G)$. If $G$ is $f$-choosable, then we say $G$ is degree-truncated $k$-choosable. The degree-truncated choice number of $G$ is the minimum integer $k$ such that $G$ is degree-truncated $k$-choosable.
In the degree-truncated $k$-list colouring model, each vertex $v$ has a list of size $k$, however, if a vertex $v$ has degree $d_{G}(v)<k$, then its list size is reduced to $d_{G}(v)$. In the study of $k$-list colouring of graphs, we do not need to consider vertices of degree less than $k$, as they can always be coloured properly, no matter what colours are assigned to its neighbours. In the degree-truncated $k$-list colouring model, vertices of small degree are also critical to colourability.
If $G$ is a Gallai-tree, then $G$ is not degree-choosable, and hence not degree-truncated $k$-choosable for any $k$. The degree-truncated choice of $G$ is undefined. For any connected graph $G$, its degree-truncated choice number is at most $\Delta(G)+1$.
Degree-truncated choice number of graphs was first studied by Hutchinson [3]. A classical result of Thomassen [7] in list colouring says that the choice number of every planar graph is at most 5. A natural question is the maximum degree-truncated choice number of planar graphs.
For any positive integer $k$, the complete bipartite graph $K_{2,k^{2}}$ is planar and not degree-truncated $k$-choosable. Some other planar graphs of connectivity at most 2 are easily seen to have arbitrarily large degree-truncated choice numbers. Bruce Richter asked whether every 3-connected non-complete planar graph is degree-truncated 6-choosable (cf. [3]). Motivated by this question, Hutchinson studied the degree-truncated choice number of outerplanar graphs. It was proved in [3] that if $G\neq K_{3}$ is a 2-connected maximal outerplanar graph (i.e., each inner face is a triangle), then $G$ is degree-truncated 5-choosable. If $G$ is a 2-connected bipartite outerplanar graph, then $G$ is 4-truncated degree-choosable. The results are sharp, as there are 2-connected maximal outerplanar graphs that are not degree-truncated 4-choosable, and 2-connected bipartite outerplanar graphs that are not degree-truncated 3-choosable. Hutchinson’s result was strengthened in [5], where it was proved that 2-connected $K_{2,4}$-minor free graphs other than cycles and complete graphs are degree-truncated 5-DP-colourable. In [1], some families of $K_{5}$-minor free graphs are shown to be degree-truncated $k$-choosable for $k=6,7,8$, by putting restrictions on the distance between components of the subgraph induced by vertices of degree less than $k$.
Recently, Zhou, Zhu, and Zhu [9] answered Richter’s question in negative and constructed a 3-connected non-complete planar graph that is not degree-truncated 7-choosable. They also proved that every 3-connected non-complete planar graph is degree-truncated $16$-DP-colourable as well as degree-truncated $16$-AT, which implies that these graphs are degree-truncated $16$-choosable, as well as online degree-truncated $16$-choosable. Before this result, it was not known whether the degree-truncated choice number of 3-connected non-complete planar graphs is bounded by a constant.
@@
There is a 3-connected non-complete planar graph that is not degree-truncated 8-choosable.
@@
Every 3-connected non-complete planar graph is degree-truncated 12-choosable.
We remark that our result takes advantage of the list assignment and does not apply to DP-colouring, on-line list colouring, or Alon-Tarsi orientation.